OpenStudy (anonymous):

Nine times a number minus three is equal to five times that same number added to seven. What is the value of the number?

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

got it right this time: 9n-3=5n+7 4n=10 n=10/4=2.5

OpenStudy (radar):

I find the wording in this problem troublesom as well as the answer. I assume that you use the rule of order of operations and come up with something like this; 9x-3=5x+7 or we might look at it like this 9(x-3)=5(x+7) semantics. But irregardless how we look at it, the answer comes out as a fraction. When the problem states it is a number (whole number ???)